A 250 mL sample of a 0.20M Cr3+ is electrolysed with a current of 96.5A. If the remaining [Cr3+] is 0.1M, the duration of process is?

Grade:10

1 Answers

Amaresh
5 years ago
0.2 M is becoming 0.1M means decrease in Cr3+ concentration is 0.1M. Number of moles = M*V= 0.1*250*10-3 =25*10-3 moles.
1mole of Cr+3 requires 3F electrons to get electrolysed. So 25*10-3 moles requires 75*10-3F.
As q=i*t
75*10-3*96500 = 96.5 * t   (1F=96500 Columbs).
So t = 75 seconds.